OpenStudy (anonymous):

how do you find an angle on the coordinate plane without a protractor or a compass

OpenStudy (amistre64):

well, since its a corrdinate plane ... do some trig with a calculator

OpenStudy (anonymous):

idk how

OpenStudy (amistre64):

for a given coord (x,y); tan(angle) = y/x and other identities are similar if the distance fromthe origin can be detemrined

OpenStudy (amistre64):

 sin(a) = y/r cos(a) = x/r tan(a) = y/x and the inverses are functions on a calculator as well. if its a simple x,y there are things called special triangles that you can put to memory for comparison

OpenStudy (anonymous):

yes my teacher told us you can use special right triangles but she never told us how

OpenStudy (amistre64):

a triangle scales .... so if you know the leg lengths ,,,, you can compare that to the specials to see what fits.

OpenStudy (amistre64):

 and a 45 should be kindergarten work